OpenStudy (anonymous):

find Particular antiderivative satisfy the given condition R'(x) = 600 -0.6x; R(0) = 0

OpenStudy (anonymous):

Integrate. R(x)=600x-0.3x^2+C. R(0)=C, so C=0.

OpenStudy (anonymous):

got you.. that's what I thought but I was not confident enough to determine if it is right

OpenStudy (anonymous):

Ah, well, there you go.

OpenStudy (anonymous):

you are fast let me post another question

OpenStudy (anonymous):

so can I show you one I did to make sure I really got it

OpenStudy (anonymous):

Sure. can you fan me again as well? :D

OpenStudy (anonymous):

dR/dt = 100/t^2 R(1) = 400 so I came upt with R= 100^t +500

OpenStudy (anonymous):

I did it earlier..I can do it again

OpenStudy (anonymous):

it won't allow me to fan you again.

OpenStudy (anonymous):

You should be able to do it once per question, I think. Anyhow, you did it correctly.

OpenStudy (anonymous):

Ok.. I will post another question then I wil fan you